An investment of $100 today is worth $121 at the end of two years if it earns an annual interest rate of 8%. How much interest is earned in the first year and how much in the second year of this investment?
A) The interest earned in year one is $10.21 and the interest earned in year two is $12.21.
B) The interest earned in year one is $10.00 and the interest earned in year two is $11.
C) The interest earned in year one is $10.21 and the interest earned in year two is $18.00.
D) There is not enough information to solve this problem.
Answer #1

Let the effective annual interest rate be x

100(1+x)2 = 121

X = 10%

Hence, interest in year one = 100*10% = $10

Interest in year two = Interest in Year one + Interest earned on interest of year 1

= 10 + 10*10%

= $11

Hence, the answer is B

B) The interest earned in year one is $10.00 and the interest earned in year two is $11.